# Sparse-PGD: A Unified Framework for Sparse Adversarial Perturbations Generation

### v3

@@ Abstract
This work studies sparse adversarial perturbations, including both unstructured and structured ones. We propose a framework based on a white-box PGD-like attack method named Sparse-PGD to effectively and efficiently generate such perturbations. Furthermore, we combine Sparse-PGD with a black-box attack to comprehensively and more reliably evaluate the models’ robustness against unstructured and structured sparse adversarial perturbations. Moreover, the efficiency of Sparse-PGD enables us to conduct adversarial training to build robust models against various sparse perturbations. Extensive experiments demonstrate that our proposed attack algorithm exhibits strong performance in different scenarios. More importantly, compared with other robust models, our adversarially trained model demonstrates state-of-the-art robustness against various sparse attacks.
1 Introduction
@@
However, constructing such adversarial perturbations is challenging as the corresponding adversarial budget is non-convex. Therefore, gradient-based methods, such as projected gradient descent (PGD) [4], usually cannot efficiently obtain a strong adversarial perturbation. In this regard, existing methods to generate sparse perturbations [12, 13, 14, 15, 16, 17] either cannot control the $l_{0}$ norm or the group $l_{0}$ norm of perturbations or have prohibitively high computational complexity, which makes them inapplicable for adversarial training to obtain robust models against sparse perturbations. The perturbations bounded by $l_{1}$ norm are the closest scenario to $l_{0}$ bounded perturbations among convex adversarial budgets defined by an $l_{p}$ norm. Nevertheless, adversarial training in this case [18, 19] still suffers from issues such as slow convergence and instability. Jiang et al. [20] demonstrates that these issues arise from non-sparse perturbations bounded by $l_{1}$ norm. In other words, $l_{1}$ adversarial budget guarantees still cannot the sparsity of the perturbations. Furthermore, existing works investigating structured sparse perturbations [21, 22, 23, 16, 17] only support generating a single adversarial patch, lacking flexibility and generality. Thus, it is necessary but challenging to develop a unified framework for both unstructured and structured sparse adversarial perturbations.
@@
$\begin{split}\mathop{\min}\limits_{\bm{\theta}}\frac{1}{N}\sum_{i=1}^{N}\mathop{\max}\limits_{\bm{\delta}_{i}}\mathcal{L}(\bm{\theta},{\bm{x}}_{i}+\bm{\delta}_{i}),\\ \mathrm{s.t.}\ ||\bm{\delta}_{i}||_{p}\leq\epsilon,\ 0\leq{\bm{x}}_{i}+\bm{\delta}_{i}\leq 1.\end{split}$ (1)
@@ 2 Related Works
3 Unstructured Sparse Adversarial Attack
In this section, we introduce Sparse-PGD (sPGD) for generating unstructured sparse perturbations. Its extension that generates structured sparse perturbations is introduced in Sec. 4. Similar to AutoAttack [39, 19], we further combine sPGD with a black-box attack to construct sparse-AutoAttack (sAA) for more comprehensive and reliable robustness evaluation.
3.1 Sparse-PGD (sPGD)
@@
Update Magnitude Tensor ${\bm{p}}$: The magnitude tensor ${\bm{p}}$ is only constrained by the input domain. In the case of images, the input is bounded between 00 and $1$. Note that the constraints on ${\bm{p}}$ are elementwise and similar to those of $l_{\infty}$ bounded perturbations. Therefore, instead of greedy or random search [13, 17], we utilize PGD in the $l_{\infty}$ case, i.e., use the sign of the gradients, to optimize ${\bm{p}}$ as demonstrated by Eq. (4) below, with $\alpha$ being the step size.
@@
Update Sparsity Mask ${\bm{m}}$: The sparsity mask ${\bm{m}}$ is binary and constrained by its $l_{0}$ norm. Directly optimizing the discrete variable ${\bm{m}}$ is challenging, so we update its continuous alternative $\widetilde{{\bm{m}}}\in\mathbb{R}^{h\times w\times 1}$ and project $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{\bm{m}}$ to obtain ${\bm{m}}$ before multiplying it with the magnitude tensor ${\bm{p}}$ to obtain the sparse perturbation $\bm{\delta}$. Specifically, $\widetilde{{\bm{m}}}$ is updated by gradient ascent. Projecting $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{{\bm{m}}}$ is to set the $\epsilon$-largest elements in $\widetilde{{\bm{m}}}$ to $1$ and the rest to 00. In addition, we adopt the sigmoid function to normalize the elements of $\widetilde{{\bm{m}}}$ before projection.
@@
${\bm{m}}$ $\longleftarrow\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$ (6)
where $\beta$ is the step size for updating the sparsity mask’s continuous alternative $\widetilde{{\bm{m}}}$, $\sigma(\cdot)$ denotes the sigmoid function. Furthermore, to prevent the magnitude of $\widetilde{{\bm{m}}}$ from becoming explosively large, we do not update $\widetilde{{\bm{m}}}$ when $||\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}||_{2}<\gamma$, which indicates that $\widetilde{{\bm{m}}}$ is located in the saturation zone of sigmoid function. The gradient $\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}$ is calculated at the point $\bm{\delta}={\bm{p}}\odot\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$, where the loss function is not always differentiable. We demonstrate how to estimate the update direction in the next part.
Backward Function: Based on Eq. (3), we can calculate the gradient of the magnitude tensor ${\bm{p}}$ by $\nabla_{\bm{p}}{\mathcal{L}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot{\bm{m}}$ and use $g_{{\bm{p}}}$ to represent this gradient for notation simplicity. At most, $\epsilon$ non-zero elements are in the mask ${\bm{m}}$, so $g_{\bm{p}}$ is sparse and has at most $\epsilon$ non-zero elements. That is to say, we update at most $\epsilon$ elements of the magnitude tensor ${\bm{p}}$ based on the gradient $g_{\bm{p}}$. Like coordinate descent, this may result in suboptimal performance since most elements of ${\bm{p}}$ are unchanged in each iterative update. To tackle this problem, we discard the projection to the binary set ${\mathcal{S}}_{{\bm{m}}}$ when calculating the gradient and use the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ to update ${\bm{p}}$. Based on Eq. (6), we have $\widetilde{g}_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}(\bm{\theta},{\bm{x}}+\bm{\delta})\odot\sigma(\widetilde{{\bm{m}}})$. The idea of the unprojected gradient is inspired by training pruned neural networks and lottery ticket hypothesis [60, 61, 62, 63]. All these methods train importance scores to prune the model parameters but update the importance scores based on the whole network instead of the pruned sub-network to prevent the sparse update, which leads to suboptimal performance.
@@
When calculating the gradient of the continuous alternative $\widetilde{{\bm{m}}}$, we have $\frac{\partial{\mathcal{L}}}{\partial\widetilde{{\bm{m}}}}=\frac{\partial{\mathcal{L}}(\theta,{\bm{x}}+\bm{\delta})}{\partial\bm{\delta}}\odot{\bm{p}}\odot\frac{\partial\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))}{\partial\widetilde{{\bm{m}}}}$. Since the projection to the set ${\mathcal{S}}_{{\bm{m}}}$ is not always differentiable, we discard the projection operator and use the approximation $\frac{\partial\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))}{\partial\widetilde{{\bm{m}}}}\simeq\sigma^{\prime}(\widetilde{{\bm{m}}})$ to calculate the gradient.
@@
7:        $g_{{\bm{p}}}=\nabla_{\bm{\delta}}{\mathcal{L}}\odot\sigma(\widetilde{{\bm{m}}})$       $\bm{\delta}={\bm{p}}\odot{\bm{m}}$}
@@
11:     $g_{\widetilde{{\bm{m}}}}=\nabla_{\bm{\delta}}{\mathcal{L}}\odot{\bm{p}}\odot\sigma^{\prime}(\widetilde{{\bm{m}}})$
@@
13:     ${\bm{d}}=g_{\widetilde{{\bm{m}}}}/\|g_{\widetilde{{\bm{m}}}}\|_{2}$ if $||g_{\widetilde{{\bm{m}}}}||_{2}\geq\gamma$ else 00
@@
15:     ${\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}))$
@@
19:     if $||{\bm{m}}-{\bm{m}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
@@
3.2 Sparse-AutoAttack (sAA)
AutoAttack (AA) [39] is an ensemble of four diverse attacks for a standardized parameters-free and reliable evaluation of robustness against $l_{\infty}$ and $l_{2}$ attacks. Croce and Hein [19] extends AutoAttack to $l_{1}$ bounded perturbations. In this work, we propose sparse-AutoAttack (sAA), which is also a parameter-free ensemble of both black-box and white-box attacks for comprehensive robustness evaluation against $l_{0}$ bounded perturbations. It can be used in a plug-and-play manner. However, different from the $l_{\infty}$, $l_{2}$ and $l_{1}$ cases, the adaptive step size, momentum and difference of logits ratio (DLR) loss function do not improve the performance in the $l_{0}$ case, so they are not adopted in sAA. In addition, compared with targeted attacks, sPGD turns out stronger when using a larger query budget in the untargeted settings given the same total number of back-propagations. As a result, we only include the untargeted sPGD with cross-entropy loss and constant step sizes in sAA. Specifically, we run sPGD twice for two different backward functions: one denoted as sPGDproj uses the sparse gradient $g_{{\bm{p}}}$, and the other denoted as sPGDunproj uses the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ as described in Section 3.1. As for the black-box attack, we adopt the strong black-box attack Sparse-RS [17], which can generate $l_{0}$ bounded perturbations. We run each version of sPGD and Sparse-RS for $10000$ iterations, respectively. We use cascade evaluation to improve the efficiency. Concretely, suppose we find a successful adversarial perturbation by one attack for one instance. Then, we will consider the model non-robust in this instance and the same instance will not be further evaluated by other attacks. Based on the efficiency and the attack success rate, the attacks in sAA are sorted in the order of sPGDunproj, sPGDproj and Sparse-RS.
4 Structured Sparse Adversarial Attack
@@
4.1 Formulation of Structured Sparsity
@@
$=F(\{\pi_{1},...,\pi_{d^{\prime}}\})\ \ \mathrm{where}\ x_{\pi_{d^{\prime}}}\neq 0\ \mathrm{and}\ x_{\pi_{d^{\prime}+1}}=0.$
@@
Figure 1: The ratio between the group $l_{0}$ norm and the approximated group $l_{0}$ norm. The x-axis is the approximated group $l_{0}$ norm, and the y-axis is the ratio. We plot the ratios of random and adversarial $3\times 3$ patch perturbations, where the adversarial perturbations are generated on vanilla ResNet-18 trained on CIFAR-10. The solid line and shadow denote the mean value and standard deviation, respectively.
Theorem 4.1 indicates that the number of perturbed groups determined by ${\bm{v}}$ can be larger than the minimal number of groups to cover the perturbed features ${\bm{x}}$, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ can be larger than $\Omega_{0}({\bm{x}})$. The gap between $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ and $\Omega_{0}({\bm{x}})$ stem from the potential overlap among groups in ${\mathcal{G}}$. However, we demonstrate that $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ is quite close to $\Omega_{0}({\bm{\delta}})$ in practice. To corroborate this, we calculate the ratio between the group $l_{0}$ norm and approximated group $l_{0}$ norm, i.e., $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ of both random and adversarial perturbations under different sparsity level $\epsilon$ defined by $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})\leq\epsilon$. As illustrated in Figure 1, when applying $3\times 3$ patch perturbations, we can observe that the ratio $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ decreases as $\epsilon$ increases. Specifically, when $\epsilon$ is smaller than $5$, the ratios of both random and adversarial perturbations are larger than $0.95$. Notably, the $l_{0}$ norm of $5$ non-overlapping $3\times 3$ patches reaches $45$, which is very large for a sparse perturbation used in practice, so we can conclude based on the simulation results that $\Omega_{0}^{\prime}$ is a good approximation of $\Omega_{0}$ in the context of the structured sparse attack.
4.2 sPGD for Structured Sparse Perturbations
@@
Based on the definition of structured sparsity and $\Omega^{\prime}_{0}$ defined in Eq. (9), if a specific group $G_{i}$ is chosen to be perturbed, then we let $v_{i}=1$, otherwise $v_{i}$ will be set 00. In this regard, finding the optimal ${\bm{m}}$ s.t. $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\leq\epsilon$ is equivalent to finding the optimal group mask ${\bm{v}}$ s.t. $\|{\bm{v}}\|_{0}\leq\epsilon$, which can be resolved by sPGD. Similar to the unstructured cases, we optimize the continuous alternative $\widetilde{{\bm{v}}}$ rather than optimizing a binary ${\bm{v}}$ directly. In the following, we will elaborate on the mapping from ${\bm{v}}$ to ${\bm{m}}$ using different examples.
@@
3:  ${\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}))$ where ${\mathcal{S}}_{{\bm{v}}}=\{{\bm{v}}|\|{\bm{v}}\|_{0}\leq\epsilon\}$.
4:  ${\bm{m}}=\min\left(\mathrm{TConv}({\bm{v}},{\bm{k}},s),1\right)$
5:  for $i=0,1,...,T-1$ do
@@
12:     $g_{\widetilde{{\bm{v}}}}=\mathrm{Conv}(\nabla_{{\bm{m}}}{\mathcal{L}},{\bm{k}},s)\odot\sigma^{\prime}(\widetilde{{\bm{v}}})$
@@
14:     ${\bm{d}}=g_{\widetilde{{\bm{v}}}}/\|g_{\widetilde{{\bm{v}}}}\|_{2}$ if $||g_{\widetilde{{\bm{v}}}}||_{2}\geq\gamma$ else 00
@@
16:     ${\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}))$
@@
21:     if $||{\bm{v}}-{\bm{v}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
@@
4.3 sAA for Structured Sparse Perturbations
Similar to Section 3.2, we also combine sPGDunproj, sPGDproj and Sparse-RS to build Sparse-AutoAttack (sAA) for reliable evaluation of robustness against structured sparse perturbations. We extended the original version of Sparse-RS so that it can generate structured sparse perturbations with any customized patterns. Specifically, we just apply a pattern mask to the patches generated by Sparse-RS.
5 Adversarial Training

### v4

@@ Abstract
This work studies sparse adversarial perturbations, including both unstructured and structured ones. We propose a framework based on a white-box PGD-like attack method named Sparse-PGD to effectively and efficiently generate such perturbations. Furthermore, we combine Sparse-PGD with a black-box attack to comprehensively and more reliably evaluate the models’ robustness against unstructured and structured sparse adversarial perturbations. Moreover, the efficiency of Sparse-PGD enables us to conduct adversarial training to build robust models against various sparse perturbations. Extensive experiments demonstrate that our proposed attack algorithm exhibits strong performance in different scenarios. More importantly, compared with other robust models, our adversarially trained model demonstrates state-of-the-art robustness against various sparse attacks. Codes are available at https://github.com/CityU-MLO/sPGD.
I Introduction
@@
However, constructing such adversarial perturbations is challenging as the corresponding adversarial budget is non-convex. Therefore, gradient-based methods, such as projected gradient descent (PGD) [4], usually cannot efficiently obtain a strong adversarial perturbation. In this regard, existing methods to generate sparse perturbations [12, 13, 14, 15, 16, 17] either cannot control the $l_{0}$ norm or the group $l_{0}$ norm of perturbations or have prohibitively high computational complexity, which makes them inapplicable for adversarial training to obtain robust models against sparse perturbations. The perturbations bounded by $l_{1}$ norm are the closest scenario to $l_{0}$ bounded perturbations among convex adversarial budgets defined by an $l_{p}$ norm. Nevertheless, adversarial training in this case [18, 19] still suffers from issues such as slow convergence and instability. Jiang et al. [20] demonstrates that these issues arise from non-sparse perturbations bounded by $l_{1}$ norm. In other words, $l_{1}$ adversarial budget cannot guarantee the sparsity of the perturbations. Furthermore, existing works investigating structured sparse perturbations [21, 22, 23, 16, 17] only support generating a single adversarial patch, lacking flexibility and generality. Thus, it is necessary but challenging to develop a unified framework for both unstructured and structured sparse adversarial perturbations.
@@
[…]
@@ II Related Works
III Unstructured Sparse Adversarial Attack
In this section, we introduce Sparse-PGD (sPGD) for generating unstructured sparse perturbations. Its extension that generates structured sparse perturbations is introduced in Sec. IV. Similar to AutoAttack [39, 19], we further combine sPGD with a black-box attack to construct sparse-AutoAttack (sAA) for more comprehensive and reliable robustness evaluation.
III-A Sparse-PGD (sPGD)
@@
Update Magnitude Tensor $\bm{p}$}: The magnitude tensor $\bm{p}$} is only constrained by the input domain. In the case of images, the input is bounded between 0 and $1$. Note that the constraints on $\bm{p}$} are elementwise and similar to those of $l_{\infty}$ bounded perturbations. Therefore, instead of greedy or random search [13, 17], we utilize PGD in the $l_{\infty}$ case, i.e., use the sign of the gradients, to optimize $\bm{p}$} as demonstrated by Eq. (4) below, with $\alpha$ being the step size.
@@
Update Sparsity Mask $\bm{m}$}: The sparsity mask $\bm{m}$} is binary and constrained by its $l_{0}$ norm. Directly optimizing the discrete variable $\bm{m}$} is challenging, so we update its continuous alternative $\widetilde{{\bm{m}}}\in\mathbb{R}^{h\times w\times 1}$ and project $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{\bm{m}}$ to obtain $\bm{m}$} before multiplying it with the magnitude tensor $\bm{p}$} to obtain the sparse perturbation $\bm{\delta}$. Specifically, $\widetilde{{\bm{m}}}$ is updated by gradient ascent. Projecting $\widetilde{{\bm{m}}}$ to the feasible set ${\mathcal{S}}_{{\bm{m}}}$ is to set the $\epsilon$-largest elements in $\widetilde{{\bm{m}}}$ to $1$ and the rest to 0. In addition, we adopt the sigmoid function to normalize the elements of $\widetilde{{\bm{m}}}$ before projection.
@@
${\bm{m}}$ $)\displaystyle\longleftarrow\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$)) (6)
where $\beta$ is the step size for updating the sparsity mask’s continuous alternative $\widetilde{{\bm{m}}}$, $\sigma(\cdot)$ denotes the sigmoid function. Furthermore, to prevent the magnitude of $\widetilde{{\bm{m}}}$ from becoming explosively large, we do not update $\widetilde{{\bm{m}}}$ when $\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}||_{2}<\gamma$, which indicates that $\widetilde{{\bm{m}}}$ is located in the saturation zone of sigmoid function. The gradient $\nabla_{\widetilde{{\bm{m}}}}{\mathcal{L}}$ is calculated at the point $)\bm{\delta}={\bm{p}}\odot\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$)), where the loss function is not always differentiable. We demonstrate how to estimate the update direction in the next part.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
13:  ${\bm{d}}=g_{\widetilde{{\bm{m}}}}/\|g_{\widetilde{{\bm{m}}}}\|_{2}$ if $g_{\widetilde{{\bm{m}}}}||_{2}\geq\gamma$ else 0
@@
15:  $){\bm{m}}=\Pi_{{\mathcal{S}}_{{\bm{m}}}}(\sigma(\widetilde{{\bm{m}}}$))
@@
19:  if ${\bm{m}}-{\bm{m}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
@@
III-B Sparse-AutoAttack (sAA)
AutoAttack (AA) [39] is an ensemble of four diverse attacks for a standardized parameters-free and reliable evaluation of robustness against $l_{\infty}$ and $l_{2}$ attacks. Croce and Hein [19] extends AutoAttack to $l_{1}$ bounded perturbations. In this work, we propose sparse-AutoAttack (sAA), which is also a parameter-free ensemble of both black-box and white-box attacks for comprehensive robustness evaluation against $l_{0}$ bounded perturbations. It can be used in a plug-and-play manner. However, different from the $l_{\infty}$, $l_{2}$ and $l_{1}$ cases, the adaptive step size, momentum and difference of logits ratio (DLR) loss function do not improve the performance in the $l_{0}$ case, so they are not adopted in sAA. In addition, compared with targeted attacks, sPGD turns out stronger when using a larger query budget in the untargeted settings given the same total number of back-propagations. As a result, we only include the untargeted sPGD with cross-entropy loss and constant step sizes in sAA. Specifically, we run sPGD twice for two different backward functions: one denoted as sPGDproj uses the sparse gradient $g_{{\bm{p}}}$, and the other denoted as sPGDunproj uses the unprojected gradient $\widetilde{g}_{{\bm{p}}}$ as described in Section III-A. As for the black-box attack, we adopt the strong black-box attack Sparse-RS [17], which can generate $l_{0}$ bounded perturbations. We run each version of sPGD and Sparse-RS for $10000$ iterations, respectively. We use cascade evaluation to improve the efficiency. Concretely, suppose we find a successful adversarial perturbation by one attack for one instance. Then, we will consider the model non-robust in this instance and the same instance will not be further evaluated by other attacks. Based on the efficiency and the attack success rate, the attacks in sAA are sorted in the order of sPGDunproj, sPGDproj and Sparse-RS.
IV Structured Sparse Adversarial Attack
@@
IV-A Formulation of Structured Sparsity
@@
$=F(\{\pi_{1},.,\pi_{d^{\prime}}\})\ \ \mathrm{where}\ x_{\pi_{d^{\prime}}}\neq 0\ \mathrm{and}\ x_{\pi_{d^{\prime}+1}}=0.$
@@
Theorem IV.1 indicates that the number of perturbed groups determined by $\bm{v}$} can be larger than the minimal number of groups to cover the perturbed features $\bm{x}$}, i.e., $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ can be larger than $\Omega_{0}({\bm{x}})$. The gap between $\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})$ and $\Omega_{0}({\bm{x}})$ stem from the potential overlap among groups in $\mathcal{G}$}. For instance, if we generate $1\times 3$ perturbations, for a $\bm{x}$} with the shape of $1\times 5$ and all of its entities are perturbed, the corresponding $\Omega_{0}({\bm{x}})=2$, while the possible $\|{\bm{v}}\|_{0}$ can be $2$ or $3$, i.e., ${\bm{v}}=[1,0,1]$ or $[1,1,1]$. Thus, we have $\Omega_{0}({\bm{x}})\leq\Omega_{0}^{\prime}({\bm{x}},{\bm{v}})\leq\|{\bm{v}}\|_{0}$. However, we demonstrate that $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ is quite close to $\Omega_{0}({\bm{\delta}})$ in practice. To corroborate this, we calculate the ratio between the group $l_{0}$ norm and approximated group $l_{0}$ norm, i.e., $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ of perturbations under different sparsity levels $\epsilon$ defined by $\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})\leq\epsilon$. As illustrated in Figure 1, when applying patch perturbations, we can observe that the ratio $\Omega_{0}({\bm{\delta}})/\Omega_{0}^{\prime}({\bm{\delta}},{\bm{v}})$ increases as the ratio between the maximum number of perturbed pixels and the total number of pixels decreases, because the probability of overlapping accordingly declines. Specifically, when $\epsilon$ is smaller than $5$, the ratios of perturbations on both CIFAR-10 and ImageNet-100 are larger than $0.95$. Notably, on ImageNet-100, the $l_{0}$ norm of $5$ non-overlapping $10\times 10$ patches reaches $500$, which is very large for a sparse perturbation used in practice, so we can conclude based on the simulation results that $\Omega_{0}^{\prime}$ is a good approximation of $\Omega_{0}$ across datasets with different resolutions.
Figure 1: The ratio between the group $l_{0}$ norm and the approximated group $l_{0}$ norm. The x-axis is the approximated group $l_{0}$ norm ranging from 1 to 25, and the y-axis is the ratio. We plot the ratios of $3\times 3$ patch perturbations on CIFAR-10 ($32\times 32$) and $10\times 10$ patch perturbations on ImageNet-100 ($224\times 224$), respectively. The results are calculated on 100 samples. The solid line and shadow denote the mean value and standard deviation, respectively.
IV-B sPGD for Structured Sparse Perturbations
@@
Based on the definition of structured sparsity and $\Omega^{\prime}_{0}$ defined in Eq. (9), if a specific group $G_{i}$ is chosen to be perturbed, then we let $v_{i}=1$, otherwise $v_{i}$ will be set 0. In this regard, finding the optimal $\bm{m}$} s.t. $\Omega^{\prime}_{0}({\bm{m}},{\bm{v}})\leq\epsilon$ is equivalent to finding the optimal group mask $\bm{v}$} s.t. $\|{\bm{v}}\|_{0}\leq\epsilon$, which can be resolved by sPGD. Similar to the unstructured cases, we optimize the continuous alternative $\widetilde{{\bm{v}}}$ rather than optimizing a binary $\bm{v}$} directly. In the following, we will elaborate on the mapping from $\bm{v}$} to $\bm{m}$} using different examples.
@@
[…]
5:for i=0,1,,T1i=0,1,...,T-1 do
@@
[…]
@@
14:  ${\bm{d}}=g_{\widetilde{{\bm{v}}}}/\|g_{\widetilde{{\bm{v}}}}\|_{2}$ if $g_{\widetilde{{\bm{v}}}}||_{2}\geq\gamma$ else 0
@@
16:  $){\bm{v}}=\Pi_{{\mathcal{S}}_{{\bm{v}}}}(\sigma(\widetilde{{\bm{v}}}$))
@@
21:  if ${\bm{v}}-{\bm{v}}_{old}||_{0}\leq 0$ for $t$ consecutive iters then
@@
IV-C sAA for Structured Sparse Perturbations
Similar to Section III-B, we also combine sPGDunproj, sPGDproj and Sparse-RS to build Sparse-AutoAttack (sAA) for reliable evaluation of robustness against structured sparse perturbations. We extended the original version of Sparse-RS so that it can generate structured sparse perturbations with any customized patterns. Specifically, we just apply a pattern mask to the patches generated by Sparse-RS.
V Adversarial Training